# ID-Booth: Identity-consistent Face Generation with Diffusion Models

### v3

@@ I Introduction
Deep learning models are nowadays utilized as backbones in a variety of recognition systems [1]. These models typically require sufficiently diverse and large-enough training datasets to achieve competitive performance. However, obtaining suitable datasets can be difficult in the field of biometrics, due to copyright, consent and privacy issues [22, 29]. Given the recent advancements in generative models, researchers are increasingly exploring the use of synthetic data to address the data needs of contemporary deep learning models. This is especially true in the area of face recognition, where synthetic data may be used to train face recognition models, augment existing datasets and enrich the variations present in real-world datasets [6].
@@
Figure 1: Samples generated with the proposed ID-Booth framework. The framework enables fine-tuning of pretrained diffusion models for generating diverse identity-consistent images based on images gathered in a constrained setting with the consent of subjects.
@@
State-of-the-art generative models are currently dominated by diffusion-based techniques, which offer unparalleled synthesis capabilities in terms of quality and diversity of the generated data, while enabling synthesis guided by text prompts [40]. Recently, diffusion models have also been utilized to produce datasets suitable for face recognition tasks, i.e., containing images of multiple identities with multiple samples each. To this end, approaches rely on identity-conditioning [34, 51, 52] and fine-tuning [42, 35] of pretrained diffusion models. Nevertheless, most solutions focus mainly on image reconstruction during training, resulting in poor consistency between the desired identities and the generated ones. To address this issue, PortraitBooth [35] recently extended the fine-tuning DreamBooth [42] method with an identity-based training objective. However, the proposed solution only considers the identity similarity of input samples and the generated samples during training. In turn, it tends to overfit on input identity features, including undesired characteristics, e.g., the pose, age, hair, accessories, thus lowering the diversity of generated images.
@@
In this paper, we present a solution for the outlined issues, in the form of a new generative framework, called ID-Booth. The proposed framework entails three main components, including $(i)$ a denoising network that produces data based on input noise, $(ii)$ a Variational Auto-Encoder (VAE) that maps images to and from a more efficient latent space on which the denoising network operates, and $(iii)$ a text encoder that enables prompt-based conditioning of the denoising network. The proposed framework utilizes a novel triplet identity objective, which considers both positive and negative identity samples during training, to facilitate the generation of identity-consistent images while retaining the synthesis capabilities of pretrained models. Throughout the experiments, we explore the suitability of ID-Booth for addressing privacy concerns by generating diverse synthetic in-the-wild images of identities from the Tufts Face Database [33], which contains images gathered in a constrained setting with the consent of subjects, as shown in Figure 1. We perform fine-tuning of a state-of-the-art diffusion model conditioned on diverse prompts and compare synthesis results with DreamBooth [42] and PortraitBooth [35] in terms of image quality, fidelity and diversity as well as intra-identity consistency and inter-identity separability. Furthermore, we investigate the real-world utility of the produced synthetic samples for augmenting existing datasets to train modern face recognition models in a privacy-preserving manner. We demonstrate that our fine-tuning framework enables the generation of more diverse synthetic samples with better intra-identity consistency and inter-identity separability. As showcased by improved recognition performance, across five real-world verification benchmarks, this makes our approach more suitable for augmenting small-scale training datasets than DreamBooth [42] or PortraitBooth [35]. Overall, the paper makes the following contributions:
@@ III-A The ID-Booth framework
The proposed diffusion-based ID-Booth framework consists of three primary components, as depicted in Figure 2. This includes $(i)$ the denoising network, responsible for enabling data generation through diffusion, $(ii)$ the Variational Auto-Encoder (VAE) that is used to map images to and from a more efficient latent space, and $(iii)$ the text encoder, which enables prompt-based control of the generation process. Fine-tuning of pretrained diffusion models is then achieved with three training objectives, $(i)$ the conventional reconstruction loss on a small set of input samples, $(ii)$ the prior preservation loss, focused on combating overfitting through the reconstruction of images generated by the model before training, and $(iii)$ the triplet identity loss, which utilizes a pretrained face recognition model to guide the diffusion model toward better similarity of generated identities and desired identities in input samples rather than random identities in prior images. Details of each component and objective are provided below.
@@ III-B Training objectives of ID-Booth
Afterward, we can decode the estimated denoised latent $\hat{z}_{0}$ to the estimated input image with $\hat{x}_{0}=\mathcal{D}(\hat{z}_{0})$. Next, we extract the facial region with a face detection model for both the estimated and the input training image, denoted as $\hat{x}^{f}_{0}$ and $x^{f}_{0}$ respectively. If the facial region exist, we obtain the identity feature representations of each image with a pretrained face recognition model $\varphi$.
@@ IV Experiments and Results
Implementation details. We evaluate the suitability of our framework on the state-of-the-art diffusion model Stable Diffusion 2.1 (SD-2.1) [40], which is capable of generating high-quality and diverse images at a resolution of $512\times 512$ through $1000$ denoising timesteps, specified by the discrete denoising scheduler with $\beta_{end}=0.012$, $beta_{start}=8.5\times 10^{-4}$ [18]. We fine-tune the SD-2.1 model on images of each identity in the Tufts Face Database (TFD) [33]. To this end, we utilize the training objectives specified by either DreamBooth [42], focused primarily on image reconstruction, PortraitBooth [42], which includes a simple two-point identity objective, or by our proposed ID-Booth framework, that balances identity consistency and image diversity through $\mathcal{L}_{Total}$. The identity objectives are based on features of a ResNet-100 face recognition model trained with the ArcFace loss [9] from face regions of noisy samples detected with MTCNN [53]. The detection of faces also acts as the decision factor for when identity-based training objectives are applied. To minimize the effect on the synthesis capabilities of the pretrained model, we utilize the Low-Rank Adaptation (LoRA) [19] method, which freezes the diffusion model but introduces new trainable layers instead. Specifically, we add two linear layers of rank $4$ to each cross-attention block, initialized with a Gaussian distribution. We also generate $200$ images with the initial SD-2.1 model and the prompt photo of a person, which are used for preservation of prior concepts through $\mathcal{L}_{PR}$ [42]. We then perform fine-tuning with images of a desired identity and the prompt photo of [ID] person, where [ID] represents a rare text token that will be tied to the new identity, in our case sks, following existing works [42]. We utilize an initial learning rate of $10^{-4}$ and the AdamW optimizer [27] with $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$ and a weight decay of $0.01$, along with the half-precision floating point format to lower VRAM usage. The fine-tuning process is stopped after $32$ epoch (i.e., $6400$ steps), based on our initial observations and existing works [42, 35].
Data generation.  Each fine-tuned SD model is used to generate two synthetic datasets, one with $21$ images per identity, as is the case in TFD [33], and one with $100$ images per identity to investigate the scalability of our approach. Data generation is performed with a guidance scale of $5.0$ and $30$ inference denoising steps with the same discrete denoising scheduler as during training. The goal is to generate diverse synthetic images of desired identities under various scenarios. To produce images that resemble real-world in-the-wild datasets [24], we utilize a prompt that defines a face image of a specific identity as well as the environment the image is taken:
@@
To produce a variety of poses we randomly select whether the image should be a portrait or a side-portrait, represented by [P]. In addition, we rely on a negative prompt to ensure the synthesis of more realistic images:
@@
TABLE I: Quantitative evaluation of quality, fidelity and diversity of synthetic images. Quality is assessed with Fréchet Distance [17] and Kernel Distance [2], while fidelity and diversity are measured through Density and Coverage [31]. Results are computed by comparing distributions of features extracted with DINOv2-ViT-L/14 [32] from synthetic images and real-world images of FFHQ [24], considering either entire images or only the face region. Vendi score [12] is used to evaluate intra-identity diversity, while CR-FIQA [3] measures face image quality of each sample, without a reference dataset.
@@
Experimental hardware. The experiments were conducted on a cluster of $4$ Nvidia A100 SXM4 40GB GPUs as well as a Desktop PC with an AMD Ryzen 7 7800X3D CPU with $128$ GB of RAM and an Nvidia RTX $4090$ GPU.
@@ IV-A Evaluation of generated images
Image quality. We begin our evaluation by assessing the overall quality of images, produced by either the proposed ID-Booth framework, or its two main competitors, DreamBooth [42] and PortraitBooth [35]. To this end, utilize the Fréchet Distance [17] and Kernel Distance [2] computed between synthetic features and features of the real-world FFHQ [24] dataset extracted with DINOv2-ViT-L/14 [32]. From results reported in Table I and samples in Figures 1 and 3, we can discern that with the SD-2.1 model and our defined prompts we can generate images that better match the quality of in-the-wild FFHQ [24] images than the real-world images of TFD [33], which were gathered in a constrained environment. Comparing results of the different fine-tuning methods, we see that our ID-Booth framework achieves the best quality results, scoring closest to the non fine-tuned model, while enabling identity-specific generation. This is the case both when evaluating entire images or only their face regions. Interestingly, both approaches with identity-based training objectives (i.e., PortraitBooth [35] and ID-Booth) also generate images of a better quality than DreamBooth [42]. In addition, we evaluate the quality of each face region with CR-FIQA [3]. Here, however, high face quality is not necessarily as desired as having a mix of high and low quality images, which can lead to better performing recognition models. This difference can also be observed on real-world datasets, where in-the-wild images of FFHQ [24] achieve a lower mean but higher standard deviation than constrained images of TFD [33]. Similarly, compared to DreamBooth [42] and PortraitBooth [35], our ID-Booth framework achieves a higher standard deviation but lower mean of CR-FIQA [3] scores, that are closer to those of diverse samples of the non fine-tuned models.
@@ IV-B Recognition-based experiments
Identity consistency and separability. To determine the suitability of generated images for forming recognition datasets we must also examine the consistency and separability of identities in the images. To this end, we form genuine and imposter distributions either only among synthetic identities or between synthetic and real-world identities, based on the similarity of features extracted with the pretrained ArcFace-based recognition model [9]. From verification results in Table III that describe these distributions, we can discern that both PortraitBooth [35] and ID-Booth achieve better identity consistency and separability than DreamBooth [42], due to the use of identity-based training objectives. This is the case both either among synthetic or between synthetic and real identities across all verification metrics. The only exception is the FNMR1000 score in the latter scenario, which can likely be attributed to a handful of outliers, especially when considering the lower FNMR100 scores and improved FDR values. Compared to PortraitBooth [35], our ID-Booth achieves lower FNMR scores among synthetic identities along with lower FMR and FNMR scores between synthetic and real identities, indicating fewer outliers. It also achieves a notably higher FDR score in the second scenario, which in combination with above observations signifies better intra-identity consistency and inter-identity separability. This is further supported by ID-Booth samples in Figure 4, which showcase better consistency between generated and real identities or among different synthetic samples of the same identity. Overall, these results highlight crucial characteristics of ID-Booth that demonstrate its suitability for augmenting existing datasets in a privacy-preserving manner by producing identity-consistent in-the-wild images of real-world identities from the training dataset for which we have consent.
@@
Training face recognition models. Lastly, we also explore the utility of the generated data in a real-world scenario for training deep face recognition models. As part of our experiments, we utilize the produced synthetic datasets to augment the small-scale real-world Tufts Face Database (TFD) [33] with in-the-wild syntethic images of its identities, images of which were gathered with consent in a constrained laboratory setting. The augmented datasets are then used to train a ResNet-50 [15] recognition model with the AdaFace loss [25], following the procedure described in Section II. The suitability of the synthetic datasets obtained with different fine-tuning methods is then evaluated based on the performance of the trained recognition model on five state-of-the-art face verification benchmarks.
We first explore augmentation with synthetic datasets consisting of $21$ samples per identity, denoted as SD-2.1 (21), which matches the scale of TFD [33]. Results reported in Table IV reveal that this form of augmentation enables the training of drastically better performing recognition models compared to training on only real-world samples of TFD [33]. Among the fine-tuning approaches, our ID-Booth framework achieves the highest overall accuracy, with improvements over DreamBooth [42] and PortraitBooth [35] being especially evident on the AgeDB-30 [30] and CP-LFW [55] benchmarks. In total, the use ID-Booth for augmentation results in a $6.1\%$ average accuracy increase over the recognition model trained on the non-augmented dataset. In the second set of experiments, we investigate the augmentation capabilities of larger synthetic datasets consisting of $100$ samples per identity, denoted as SD-2.1 (100). From results in Table IV we can discern notable accuracy improvements across with all fine-tuning approaches, showcasing their scalability. Similar to previous results, our ID-Booth framework achieves higher overall verification accuracy than DreamBooth [42] or PortraitBooth [35], with a total average augmentation improvement of $7.9\%$. As before, the most noticeable improvement is observed on the cross-age benchmarks (i.e., AgeDB-30 [30] and CA-LFW [56]). Overall, the accuracy improvements obtained with our ID-Booth framework can be attributed to the novel triplet identity objective, which facilitates better identity consistency and higher intra-identity diversity, especially in terms of age and pose, than either DreamBooth [42] or PortraitBooth [35].

### v4

@@ I Introduction
Deep learning models are nowadays utilized as backbones in a variety of recognition systems [1]. These models typically require sufficiently large and diverse training datasets to achieve competitive performance. However, obtaining suitable datasets can be difficult in the field of biometrics, due to copyright, consent, and privacy issues [22, 29]. With recent advancements in generative models, researchers are increasingly exploring the use of synthetic data to address the data needs of contemporary deep learning models. This is especially true in face recognition, where synthetic data may be used to train models or augment existing datasets by enriching the variation present in real-world data [6].
@@
Figure 1: Samples generated with the proposed ID-Booth framework. The framework enables fine-tuning of pretrained diffusion models for generating diverse identity-consistent face images based on images gathered in a constrained setting with the consent of subjects.
@@
State-of-the-art generative models are currently dominated by diffusion-based techniques, which offer unparalleled synthesis capabilities in terms of quality and diversity of the generated data, while enabling synthesis guided by text prompts [40]. Recently, diffusion models have also been utilized to produce datasets suitable for face recognition tasks, i.e., containing images of multiple identities with multiple samples each. To this end, approaches rely on identity-conditioning [34, 51, 52] and fine-tuning [42, 35] of pretrained diffusion models. Nevertheless, most solutions focus mainly on image reconstruction during training, resulting in poor consistency between the desired identities and the generated ones. To address this issue, PortraitBooth [35] recently extended the fine-tuning DreamBooth [42] method with an identity-based training objective. However, the proposed solution only considers the identity similarity of input samples and the generated samples during training. In turn, it tends to overfit on input identity features, including undesired characteristics, e.g., the pose, age, hair, accessories, thereby reducing the diversity of generated images.
@@
In this paper, we present a solution for the outlined issues, in the form of a new generative framework, called ID-Booth. The proposed framework entails three main components, including $(i)$ a denoising network that produces data based on input noise, $(ii)$ a Variational Auto-Encoder (VAE) that maps images to and from a more efficient latent space on which the denoising network operates, and $(iii)$ a text encoder that enables prompt-based conditioning of the denoising network. The proposed framework utilizes a novel triplet identity objective, which considers both positive and negative identity samples during training, to facilitate the generation of identity-consistent images while retaining the synthesis capabilities of pretrained models. Throughout the experiments, we explore the suitability of ID-Booth for addressing privacy concerns by generating diverse synthetic in-the-wild images of identities from the Tufts Face Database [33], which contains images gathered in a constrained setting with subject consent, as shown in Figure 1. We perform fine-tuning of a state-of-the-art diffusion model conditioned on diverse prompts and compare synthesis results with DreamBooth [42] and PortraitBooth [35] in terms of image quality, fidelity and diversity as well as intra-identity consistency and inter-identity separability. Furthermore, we investigate the real-world utility of the produced synthetic samples for augmenting existing datasets to train modern face recognition models in a privacy-preserving manner. We demonstrate that our fine-tuning framework enables the generation of more diverse synthetic samples with better intra-identity consistency and inter-identity separability. As showcased by improved recognition performance, across five real-world verification benchmarks, this makes our approach more suitable for augmenting small-scale training datasets than DreamBooth [42] or PortraitBooth [35]. Overall, the paper makes the following contributions:
@@ III-A The ID-Booth framework
The proposed diffusion-based ID-Booth framework consists of three primary components, as depicted in Figure 2. This includes $(i)$ the denoising network, responsible for enabling data generation through diffusion, $(ii)$ the Variational Auto-Encoder (VAE) that maps images to and from a more efficient latent space, and $(iii)$ the text encoder, which enables prompt-based control of the generation process. Fine-tuning of pretrained diffusion models is then achieved with three training objectives, $(i)$ the conventional reconstruction loss on a small set of input samples, $(ii)$ the prior preservation loss, focused on combating overfitting via the reconstruction of images generated before fine-tuning, and $(iii)$ the triplet identity loss, which utilizes a pretrained face recognition model to guide the diffusion model toward better similarity between generated and target identities rather than random identities from prior images. Details of each component and objective are provided below.
@@ III-B Training objectives of ID-Booth
Afterward, we can decode the estimated denoised latent $\hat{z}_{0}$ to the estimated input image with $\hat{x}_{0}=\mathcal{D}(\hat{z}_{0})$. Next, we extract the facial region with a face detection model for both the estimated and the input training image, denoted as $\hat{x}^{f}_{0}$ and $x^{f}_{0}$ respectively. If the facial region exist, we obtain the identity feature representations of each image with a pretrained face recognition model $\varphi$, otherwise the objective is skipped.
@@ IV Experiments and Results
Implementation details. We evaluate the suitability of our framework on the state-of-the-art diffusion model Stable Diffusion 2.1 (SD-2.1) [40], which is capable of generating high-quality and diverse $512\times 512$ images through $1000$ denoising timesteps, specified by the discrete denoising scheduler with $\beta_{start}=8.5\times 10^{-4}$ and $\beta_{end}=0.012$ [18]. We fine-tune the SD-2.1 model on images of each identity in the Tufts Face Database (TFD) [33]. To this end, we utilize the training objectives specified by either DreamBooth [42], focused primarily on image reconstruction, PortraitBooth [42], which includes a simple two-point identity objective, or by our proposed ID-Booth framework, that balances identity consistency and image diversity. The identity objectives utilize features extracted with a pretrained ArcFace-based ResNet-100 recognition model [9] from face regions detected with MTCNN [53]. The detection of faces also acts as the decision factor for when identity-based objectives are applied. To minimize the effect on the synthesis capabilities of the pretrained model, we utilize the Low-Rank Adaptation (LoRA) [19] method, which freezes the diffusion model but introduces new trainable layers instead. Specifically, we add two linear layers of rank $4$ to each cross-attention block, initialized with a Gaussian distribution. We also generate $200$ images with the initial SD-2.1 model and the prompt photo of a person, which are used for preservation of prior concepts through $\mathcal{L}_{PR}$ [42]. We then perform fine-tuning with images of a desired identity and the prompt photo of [ID] person, where [ID] represents a rare text token that will be tied to the new identity, in our case sks [42]. We utilize an initial learning rate of $10^{-4}$ and the AdamW optimizer [27] with $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$ and a weight decay of $0.01$, along with the half-precision floating point format to lower VRAM usage. Fine-tuning is stopped after $32$ epochs (i.e., $6400$ steps), based on our initial observations and existing works [42, 35].
Data generation.  Each fine-tuned SD model is used to generate two synthetic datasets, one with $21$ images per identity, as is the case in TFD [33], and one with $100$ images per identity to investigate the scalability of our approach. Data generation is performed with a guidance scale of $5.0$ and $30$ inference denoising steps with the same discrete denoising scheduler as during training. The goal is to generate diverse synthetic images of desired identities under various scenarios. To produce images that resemble real-world in-the-wild datasets [24], we utilize a prompt that defines a face image of a specific identity as well as the environment the image is taken in:
@@
To also produce a variety of poses we randomly select whether the image is a portrait or a side-portrait, represented by [P]. In addition, we rely on the following negative prompt to obtain more realistic images:
@@
TABLE I: Quantitative evaluation of quality, fidelity and diversity of synthetic images. Quality is assessed with Fréchet Distance [17] and Kernel Distance [2], while fidelity and diversity are measured through Density and Coverage [31]. Results are computed by comparing distributions of features extracted with DINOv2-ViT-L/14 [32] from synthetic images and real-world images of FFHQ [24], considering either entire images or only the face region. Vendi score [12] is used to evaluate intra-identity diversity, while CR-FIQA [3] measures face image quality of each sample, both without a reference dataset.
@@
Experimental hardware. The experiments were conducted on a cluster of $4$ Nvidia A100 SXM4 40GB GPUs and a Desktop PC with an Nvidia RTX $4090$ GPU.
@@ IV-A Evaluation of generated images
Image quality. We begin our evaluation by assessing the overall quality of images produced by either the proposed ID-Booth framework or its two main competitors, DreamBooth [42] and PortraitBooth [35]. To this end, we utilize the Fréchet Distance [17] and Kernel Distance [2] computed between features extracted with DINOv2-ViT-L/14 [32] from synthetic images and features extracted from real-world images of the FFHQ [24] dataset. From results reported in Table I and samples in Figures 1 and 3, we can discern that with the SD-2.1 model and our defined prompts we can generate images that better match the quality of in-the-wild FFHQ [24] images than the real-world images of TFD [33], which were gathered in a constrained environment. Comparing results of the different fine-tuning methods, we see that our ID-Booth framework achieves the best quality results, scoring closest to the non fine-tuned model, while enabling identity-specific generation. This is the case both when evaluating entire images or only the cropped face regions. Interestingly, both approaches with identity-based training objectives (i.e., PortraitBooth [35] and our ID-Booth) also generate images of a better quality than DreamBooth [42].
In addition, we evaluate the quality of each face region with CR-FIQA [3]. Here, however, high face quality is not necessarily as desired as having a mix of high and low quality images, which can lead to better performing recognition models. This difference can also be observed on real-world datasets, where in-the-wild images of FFHQ [24] achieve a lower mean but higher standard deviation than constrained images of TFD [33]. Similarly, compared to DreamBooth [42] and PortraitBooth [35], our ID-Booth framework achieves a higher standard deviation but lower mean of CR-FIQA [3] scores, that are closer to those of diverse samples of the non fine-tuned models.
@@ IV-B Recognition-based experiments
Identity consistency and separability. To determine the suitability of generated images for augmenting recognition datasets we must also examine the consistency and separability of identities in the images. To this end, we form genuine and imposter distributions either only among synthetic identities or between synthetic and real-world identities, based on the similarity of features extracted with the pretrained ArcFace-based recognition model [9]. From verification results in Table III that describe these distributions, we can discern that both PortraitBooth [35] and ID-Booth achieve better identity consistency and separability than DreamBooth [42], due to the use of identity-based training objectives. This is the case both either among synthetic or between synthetic and real identities across all verification metrics. The only exception is the FNMR1000 score in the latter scenario, which can likely be attributed to a handful of outliers, especially when considering the lower FNMR100 scores and improved FDR values. Compared to PortraitBooth [35], our ID-Booth achieves lower FNMR scores among synthetic identities along with lower FMR and FNMR scores between synthetic and real identities, indicating fewer outliers. It also achieves a notably higher FDR score in the second scenario, which in combination with above observations signifies better intra-identity consistency and inter-identity separability. This is further supported by ID-Booth samples in Figure 4, which showcase better consistency between generated and real identities or among different synthetic samples of the same identity. Overall, these results highlight crucial characteristics of ID-Booth that demonstrate its suitability for augmenting existing datasets in a privacy-preserving manner by producing identity-consistent in-the-wild images of real-world identities from the training dataset collected with subject consent.
@@
Training face recognition models. Lastly, we also explore the real-world utility of the generated data for training deep face recognition models. As part of our experiments, we utilize the produced synthetic datasets to augment the small-scale real-world Tufts Face Database (TFD) [33] with in-the-wild synthetic images of its identities, whose real-world images were gathered with consent in a constrained laboratory setting. The augmented datasets are then used to train a ResNet-50 [15] recognition model with the AdaFace loss [25], following the procedure described in Section II. The suitability of the synthetic datasets obtained with different fine-tuning methods is then evaluated based on the performance of the trained recognition model on five state-of-the-art face verification benchmarks.
We first explore augmentation with synthetic datasets consisting of $21$ samples per identity, denoted as SD-2.1 (21), which matches the scale of TFD [33]. Results reported in Table IV reveal that this form of augmentation enables the training of drastically better performing recognition models compared to training on only real-world samples of TFD [33]. Among the fine-tuning approaches, our ID-Booth framework achieves the highest overall accuracy, with improvements over DreamBooth [42] and PortraitBooth [35] being especially evident on the AgeDB-30 [30] and CP-LFW [55] benchmarks. In total, the use ID-Booth for augmentation results in a $6.1\%$ average accuracy increase over the recognition model trained on the non-augmented dataset. In the second set of experiments, we investigate the augmentation capabilities of larger synthetic datasets consisting of $100$ samples per identity, denoted as SD-2.1 (100). From results in Table IV we can discern notable accuracy improvements across with all fine-tuning approaches, showcasing their scalability. Similar to previous results, our ID-Booth framework achieves higher overall verification accuracy than DreamBooth [42] or PortraitBooth [35], with a total average augmentation improvement of $7.9\%$. As before, the most noticeable improvement is observed on cross-age benchmarks (i.e., AgeDB-30 [30] and CA-LFW [56]). Overall, the accuracy improvements obtained with our ID-Booth framework can be attributed to the novel triplet identity objective, which facilitates better identity consistency and higher intra-identity diversity, especially in terms of age and pose, than either DreamBooth [42] or PortraitBooth [35].